Suppose Proctor? & Gamble? (P&G) is considering purchasing $15 million in new manufacturing equipment. If it purchases the? equipment, it will depreciate it for tax purposes on a? straight-line basis over five? years, after which the equipment will be worthless. It will also be responsible for maintenance expenses of $1.00 million per?year, paid in each of years 1 through 5. It can also lease the equipment under a true tax lease for ?$4.1 million per year for the five? years, in which case the lessor will provide necessary maintenance. Assume? P&G's tax rate is 30 % and its borrowing cost is 6.0%.

a. What is the NPV associated with leasing the equipment versus financing it with the? lease-equivalent loan?

b. What is the? break-even lease ratel—that ?is, what lease amount could? P&G pay each year and be indifferent between leasing and financing a? purchase?
Answer #1

a) Upfront Payment for P&G= $ 15 M

Depreciation= 15/3=$ 5 M

Tax Shield Per Year= 0.30 *5= $ 1.50

After Tax Maintenance Expense Per Year=1 * (1-0.30)= $ 0.70

Thus, FCF from Buying (1-5) = 1.50-0.70= $ 0.35

If it leases the After tax payments=4.1 * (1-0.30)= $ 2.87

Thus the incremental cash flows from Leasing versus buying is:

Year 0= -2.87- (-15)= 12.13

Year (1-4) = -2.87 – (0.35)= -3.22

Year 5= 0-0.35=-0.35

Discount Rate for the cash flows = 6*(1-0.30)=4.2%

Year

0

1

2

3

4

5

FCF

12.13

-3.22

-3.22

-3.22

-3.22

-0.35

Total

Discounted FCF

12.13

-3.09

-2.97

-2.85

-2.73

-0.28

0.21170

Thus, the benefit of leasing versus buying gives a benefit or NPV of $ 0.21170

b) Increase the after tax lease payments by an equal to the NPV. Therefore, after tax lease payments that will make P&G indifferent is:

4.1+0.21170= 4.31170